# DualContrast: Unsupervised Disentangling of Content and Transformations with Implicit Parameterization

### v1

@@ Abstract
Unsupervised disentanglement of content and transformation has recently drawn much research, given their efficacy in solving downstream unsupervised tasks like clustering, alignment, and shape analysis. This problem is particularly important for analyzing shape-focused real-world scientific image datasets, given their significant relevance to downstream tasks. The existing works address the problem by explicitly parameterizing the transformation factors, significantly reducing their expressiveness. Moreover, they are not applicable in cases where transformations can not be readily parametrized. An alternative to such explicit approaches is self-supervised methods with data augmentation, which implicitly disentangles transformations and content. We demonstrate that the existing self-supervised methods with data augmentation result in the poor disentanglement of content and transformations in real-world scenarios. Therefore, we developed a novel self-supervised method, DualContrast, specifically for unsupervised disentanglement of content and transformations in shape-focused image datasets. Our extensive experiments showcase the superiority of DualContrast over existing self-supervised and explicit parameterization approaches. We leveraged DualContrast to disentangle protein identities and protein conformations in cellular 3D protein images. Moreover, we also disentangled transformations in MNIST, viewpoint in the Linemod Object dataset, and human movement deformation in the Starmen dataset as transformations using DualContrast.
Conceptualization of content-transformation disentanglement (a) and comparison between existing explicit (b) and our implicit (c) unsupervised methods. Both of these methods optimize content and transformation latent spaces while reconstructing $\mathbf{x}^{\prime}$ from $\mathbf{x}$. However, explicit methods use the transformation space to infer a fixed parameter set $\gamma$ and use it to transform the feature map/image $\mathbf{x}_{c}$ generated by conditioning the decoder on content space. On the contrary, our implicit method conditions the decoder on both content and transformation to generate image $\mathbf{x}^{\prime}$ and does not restrict the transformation space to a fixed set of parameters.
@@ 1 Introduction
Figure 1: Our proposed contrastive learning-based unsupervised content-transformation disentanglement pipeline. (a) The variational inference of content and transformation factors with $L_{\text{VAE}}$. (b,c) The contrastive strategy to create positive and negative pairs w.r.t. content (b) and transformation factors (c), respectively. DualContrast combines the three parts and is trained end-to-end.
@@
In many real-world image datasets, particularly in scientific imaging domains, the object shapes being visualized may undergo multiple transformations [1, 2]. These shape-focused image datasets can thus be regarded as samples generated from two independent factors, one representing the semantic attribute termed content and the other representing the transformations. Content is regarded as the form of an object being visualized in the image [1] that is unaltered after applying any nuisance transformation. On the other hand, transformation dictates the specific state or realization of that form of the object. Taking proteins inside the cells as an example, changing the transformation factor changes the conformation of the protein where the protein identity (i.e., protein composition) does not change (Figure DualContrast: Unsupervised Disentangling of Content and Transformations with Implicit Parameterization). On the other hand, changing the content factor is analogous to changing the proteins from one identity to another through compositional change. Such a phenomenon holds for many real-world shape-focused datasets. Disentangling content and transformation factors of images can significantly facilitate several downstream tasks, including clustering [3], alignment [4, 2], zero-shot transfer [5], bias removal [6, 7], cross-domain retrieval [8, 9], shape analysis [10, 11, 4, 2, 3], etc. [11].
@@
A number of works have been done to directly or indirectly solve the content-transformation disentangling problem [12, 8, 10, 1, 4, 2], whereas many of them use annotated factors for training. However, such annotation is hard to obtain, and thus, unsupervised disentanglement methods are desired. Only a few works [1, 2, 4] have addressed the problem in a completely unsupervised setting. Among them, VITAE [1] performed explicit parameterization of transformation factors as diffeomorphic transformations and achieved noteworthy disentanglement of content and transformation in several image datasets. SpatialVAE [2] and Harmony [4] focus on representing known transformation types as transformation factors. SpatialVAE [2] performed disentanglement of two-dimensional translation and rotation, whereas Harmony [4] demonstrated disentanglement of a broader set of transformations that can be explicitly parameterized.
@@
Though these unsupervised explicit parameterization methods have demonstrated several successes, they have several major limitations in general: (1) These method uses the transformation factor to generate only a few parameterized transformations. In reality, many transformations do not have a well-defined parameterized form,   e.g., protein conformation changes, viewpoint change (LineMod), human deformation during movement (Starmen), etc. These explicit methods can not disentangle such transformation by design. (2) They use Spatial Transformer Networks (STN) [13] for inferring the transformation factor, which requires a continuous parameterization of the transformation to be disentangled. The continuity of many transformations (e.g., SO(3) rotation) in neural networks is often a concern [14]. Moreover, (3) these methods often infer spurious contents during disentanglement in real-world datasets(Fig. 2). Due to these limitations, an unsupervised content-transformation disentangling method without explicit transformation representation is much desired.
@@
Consequently, in this work, we focused on the more generalized setting of unsupervised content-transformation disentanglement without any explicit parameterization of the transformation factor. Given the theoretical study by Von et al. [15] showing the implicit disentanglement ability of standard Contrastive Learning (CL) methods, e.g., SimCLR [16], to disentangle content from style, a plausible solution could be using SimCLR with geometric data augmentation for our implicit content-transformation disentanglement task. However, these popular CL methods [16] have rarely achieved disentanglement in practice [6], and we also found that SimCLR with geometric data augmentation performs poorly in our shape-focused image datasets of interest (Table 1).
@@
To this end, we develop a novel contrastive learning method for implicit content-transformation disentanglement, termed “DualContrast”. Similar to previous content-transformation disentangling methods [17, 2, 1, 4], it is a deep generative model based on a variational autoencoder (VAE) architecture. However, unlike previous methods, it does not require labels or explicit parameterization of any latent factor. Instead, it uses a novel contrastive strategy that creates “positive” and “negative” contrastive pairs w.r.t. both content and transformation latent factor. We create negative pairs w.r.t. transformation factor and positive pairs w.r.t. content factor by simply rotating the samples. In addition, we randomly permute the samples in the batch to create a negative pair for a batch of samples w.r.t. content latent factor. Without prior knowledge, creating positive pairs w.r.t. the transformation latent factor from the input samples is not possible. To this end, we generate samples from an identical distribution of the transformation representations in the embedding space and treat them as positive pairs for the transformation latent factor. Upon creating both “positive” and “negative” pairs w.r.t. both factors, we penalize similarity between negative pairs and distance between positive pairs for each factor. In DualContrast, we combine these contrastive losses with VAE reconstruction loss and train the model in an end-to-end manner.
@@
With such a strategy, by just performing simple rotation to create contrastive pairs, DualContrast disentangles a much wider range of more complicated transformations, e.g., human deformation in starmen, viewpoint in Linemod, the writing style in MNIST, etc. With DualContrast, we, for the first time, solved the scientific problem of disentangling protein identities from protein conformations in simulated 3D protein images (Figure 6) inside the cell, known as cellular cryo-ET subtomograms. DualContrast could successfully disentangle the protein identities as content and protein conformations as transformation factors in simulated subtomograms, which is not achievable with previous methods. Overall, our qualitative and quantitative experimental results show the efficacy of DualContrast in isolating content from transformation compared to existing relevant methods.
@@
  • We work on a challenging setting of learning unsupervised disentangled representations of content and transformation in shape-focused images by removing the requirement of explicit parameterization.
  • We develop a novel contrastive generative modeling strategy that creates positive-negative pairs for both content and transformation latent factors. We introduced a novel way of designing contrastive pairs for transformation factors.
  • We show that DualContrast is highly effective in disentangling a wide range of transformations from the content in various shape-focused image datasets by only using simple rotation for creating contrastive pairs.
  • Leveraging DualContrast, we, for the first time, disentangled protein identities (i.e.composition) from protein conformations in simulated 3D cellular cryo-ET subtomograms, as a proof of principle.
@@ 2 Related Works
Disentangled Representation Learning: Our method can be regarded as a disentangled representation learning method that disentangles generative factors of variation as content-specific and transformation-specific factors. There exist many generic disentangled representation learning approaches [18, 17, 19, 20, 21, 22] where the factors of variations are interpreted by traversing through the learned factors and reconstructing samples. However, these methods do not perform very well in specifically disentangling content and transformation factors [1]. To this end, specialized methods that deal with such two factors have been introduced recently. Many of them [23, 12, 8] are partially or weakly supervised and use labels associated with one of the latent factors to train their models. However, these methods are not applicable when such supervision is unavailable beforehand, which is often the case for many real-life datasets. There are a few [1, 4, 2, 15] approaches that deal with the unsupervised disentanglement of content and transformation. However, these methods perform explicit parameterization of the transformation factor as affine or some specific parameterized transformation. In contrast, our method does not impose any explicit parameterization on the transformation latent factor and does not face the issues of the explicit methods (Section 1).
@@
Contrastive Learning based Disentanglement: Contrastive learning is the primary building block of our method. Contrastive Learning with data augmentation with existing approaches, e.g., InfoNCE loss-based SimCLR [16], etc., has been previously used by [15, 8] to isolate content from style. Though in several scenarios, style in shape focused images can be referred to as transformations, these works do not specifically focus on shape-focused real-world images. Moreover, our experiments demonstrate that SimCLR with rotation augmentation in a standard encoder-decoder framework leads to very poor disentanglement of content and transformation (Table 1). Moreover, [15] only creates positive pairs w.r.t content and negative pair w.r.t. style with data augmentation in SimCLR. We have observed this scheme does not work well in our scenario. Unlike these works, our approach uses a novel strategy to create “negative" and “positive" pairs for both content and transformation of latent factors without any InfoNCE loss.
@@
Unsupervised Content-style disentanglement: Apart from [15], several methods [6, 24, 25, 26, 27, 28] exist that perform unsupervised content-style disentanglement, focusing on natural images. These methods focus on the controllable generation of realistic natural images, mostly using ImageNet pretrained models. Unlike these methods, our work primarily focuses on disentangling content and transformation in shape-focused image datasets. Moreover, we do not depend on any ImageNet pretrained models as our images of interest differ greatly from the natural images of ImageNet.
@@
Protein Composition-Conformation Disentanglement: One of the major contributions of our work is disentangling protein identity characterized by large compositional variability as the content factor and protein states characterized by conformational variability or subtle compositional variability as transformation factor from 3D cellular cryo-electron tomography (cryo-ET) subimages or subtomograms. Previously, Harmony [4] was used to disentangle large compositional variability in cellular subtomograms. HEMNMA-3D [29] and TomoFlow [30] were used to disentangle subtle compositional or conformational variability in subtomograms using known templates. SpatialVAE [2] disentangled subtle conformational variability from 2D cryo-EM images. There exists several works on single particle cryo-EM and cryo-ET reconstruction, e.g., cryoDRGN2 [31], cryoFIRE [3], cryoAI [32], etc. that performs amortized inference of transformation ($SO(3)\times d^{2}$) and latent space representing content. However, these works mainly focus on 2D-to-3D reconstruction instead of content-transformation disentanglement. Ours is the first work to deal with both large and subtle compositional variability in two different latent spaces for 3D cryo-ET subtomograms.
@@ 3.1 Disentangling Content and Transformation
  • Condition 1: $\forall$ $T\in\mathcal{T}$, s.t., for any $\mathbf{x}\in\mathcal{X}$, $h_{c}(T(\mathbf{x}))=h_{c}(\mathbf{x})$.
  • Condition 2: $\exists$ $T\in\mathcal{T}$, s.t., for any $\mathbf{x}^{(1)},\mathbf{x}^{(2)}\in\mathcal{X}$, $h_{z}(T(\mathbf{x}^{(1)}))=h_{z}(\mathbf{x}^{(2)})$.
@@
For content-transformation disentangling, changes in $\mathcal{C}$ must be unaffected by changes in $\mathcal{Z}$ and vice versa. In our problem setting, no labels associated with content space $C$ or transformation space $Z$ are known. The exact family of transformation $\mathcal{T}$ is also not known beforehand. Thus, we aim to simultaneously learn the family of transformations $\mathcal{T}$ and the content factors unaltered by $\mathcal{T}$.
@@ 3.2 Method Overview and Notation
Our method is built upon a variational auto-encoder (VAE) based architecture (Figure 1), where we perform inference of the content and transformation factors in the data. Consider, a sample space $\mathcal{X}$, two latent distribution spaces $\Phi_{Z}$ and $\Phi_{C}$, two latent spaces $\mathcal{Z}$ and $\mathcal{C}$. We apply an encoder network $q_{\psi}:\mathcal{X}\rightarrow\mathcal{C}\times\mathcal{Z}$ on a sample space $bx\in\mathcal{X}$ to encode content distribution $\phi_{c}\in\Phi_{c}$ and transformation distribution $\phi_{z}\in\Phi_{z}$, where $(\phi_{c},\phi_{z})=q_{\psi}(\mathbf{x})$. Latent parameters $\mathbf{c}\in\mathcal{C}$, $\mathbf{z}\in\mathcal{Z}$ are drawn from distributions $\phi_{c}$ and $\phi_{z}$ respectively. These parameters are then decoded using a decoder network $p_{\theta}:\mathcal{C}\times\mathcal{Z}\rightarrow\mathcal{X}$ to produce $x_{\text{recon}}=p_{\theta}(\mathbf{c},\mathbf{z})$.
3.3 Variational Inference of Content and Transformation Factors

### v2

@@ Abstract
Unsupervised disentanglement of content and transformation is significantly important for analyzing shape focused scientific image datasets, given their efficacy in solving downstream image-based shape-analyses tasks. The existing relevant works address the problem by explicitly parameterizing the transformation latent codes in a generative model, significantly reducing their expressiveness. Moreover, they are not applicable in cases where transformations can not be readily parametrized. An alternative to such explicit approaches is contrastive methods with data augmentation, which implicitly disentangles transformations and content. However, the existing contrastive strategies are insufficient to this end. Therefore, we developed a novel contrastive method with generative modeling, DualContrast, specifically for unsupervised disentanglement of content and transformations in shape focused image datasets. DualContrast creates positive and negative pairs for content and transformation from data and latent spaces. Our extensive experiments showcase the efficacy of DualContrast over existing self-supervised and explicit parameterization approaches. With DualContrast, we disentangled protein composition and conformations in cellular 3D protein images, which was unattainable with existing disentanglement approaches.
@@ 1 Introduction
Figure 1: (a) The concept of content-transformation disentanglement, whereas changing the content changes the protein identity, and changing transformation changes the state of the protein. Explicit methods (b) use the transformation space to infer a fixed parameter set, whereas our implicit method (c) do not restrict the transformations to a fixed set of parameters. Figure uses toy protein images for visualization, they are not used for experiments in this form.
@@
In many real-world image datasets, particularly in scientific imaging domains, the object shapes being visualized may undergo multiple transformations [1, 2]. These shape focused image datasets can thus be regarded as samples generated from two independent factors, one representing the semantic attribute termed content and the other representing the transformations. Content is regarded as the form of an object being visualized in the image [1] that is unaltered after applying any nuisance transformation. On the other hand, transformation dictates the specific state or realization of that form of the object. Taking proteins inside the cells as an example, changing the transformation factor changes the conformation of the protein where the protein identity (i.e. , protein composition) does not change (Fig. 1). On the other hand, changing the content factor is analogous to changing the proteins from one identity to another through compositional change. Such a phenomenon holds for many shape focused image datasets. Disentangling content and transformation factors of images can significantly facilitate several downstream shape analysis tasks, including shape clustering [3], alignment [4, 2], zero-shot transfer [5], bias removal [6, 7], cross-domain retrieval [8, 9], etc. [10].
@@
A number of works have been done to directly or indirectly solve the content-transformation disentangling problem [11, 8, 12, 1, 4, 2], whereas many of them use annotated factors for training. However, such annotation is hard to obtain, and thus, unsupervised disentanglement methods are desired. Only a few works [1, 2, 4] have addressed the problem in a completely unsupervised setting. Among them, [1] performed explicit parameterization of transformation codes as diffeomorphic transformations and achieved noteworthy disentanglement of content and transformation in several image datasets. [2] and [4] focus on representing known transformation types as transformation codes. [2] performed disentanglement of two-dimensional translation and rotation, whereas [4] demonstrated disentanglement of a broader set of transformations that can be explicitly parameterized.
@@
Though these unsupervised explicit parameterization methods have demonstrated several successes, they have several major limitations in general: (1) These method uses the transformation codes to generate only a few parameterized transformations. In reality, many transformations do not have a well-defined parameterized form,   e.g., protein conformation changes, viewpoint change (LineMod), etc. These explicit methods can not disentangle such transformation by design. (2) They use Spatial Transformer Networks (STN) [13] for inferring the transformation codes, which requires a continuous parameterization of the transformation to be disentangled. The continuity of many transformations (e.g. , SO(3) rotation) in neural networks is often a concern [14]. Moreover, (3) these methods infer spurious contents during disentanglement in real scenarios where transformations not predefined are present in the dataset (Fig. 3). Due to these limitations, an unsupervised content-transformation disentangling method without explicit transformation representation is much desired.
@@
Consequently, in this work, we focused on the more generalized and unique setting of unsupervised content-transformation disentanglement without any explicit parameterization of the transformation code. Previously, a theoretical study by [15] showed that standard Contrastive Learning (CL) methods, e.g., SimCLR [16], can disentangle content from style, where style can be regarded as the transformations used for data augmentation. However, these popular CL methods [16] have rarely achieved any disentanglement in practice [6], and their ability in disentangling transformations other than those used for augmentation is not well-explored. Moreover, we found that SimCLR with geometric data augmentation can not disentangle any transformation well in our shape focused image datasets of interest (Table 1).
Figure 2: Our proposed contrastive learning-based unsupervised content-transformation disentanglement pipeline. (a) The variational inference of content and transformation codes with $L_{\text{VAE}}$. (b) Contrastive pair creation strategy for content and transformation codes. The process is delineated in the bottom panel. Additional visualization available in Appendix Fig. 9. (c) Contrastive losses. In DualContrast, the contrastive pair creations and reconstruction happen simultaneously, and the encoder and decoder network are optimized with both contrastive and reconstruction losses in each iteration.
@@
To this end, we develop a novel generative contrastive learning method for implicit content-transformation disentanglement, termed “DualContrast” (Fig. 2). Unlike existing contrastive strategies, it creates “positive” and “negative” contrastive pairs w.r.t. both content and transformation latent codes. We found that several shape transformations in image datasets lie closely in the normally distributed latent space that is designed to capture in-plane rotation information in a contrastive manner. Consequently, we create negative pairs w.r.t. transformation and positive pairs w.r.t. content by simply rotating the samples. Without prior knowledge, creating positive pairs w.r.t. the transformation codes directly from the input samples is not possible. To this end, we generate samples from an identical distribution of the transformation representations in the embedding space and treat them as positive pairs for the transformation codes. Given our observation, this encourages the disentanglement of several shape transformations other than rotation in our dataset. In addition, we randomly permute the samples in the batch to create a negative pair for a batch of samples w.r.t. content. Upon creating both “positive” and “negative” pairs w.r.t. both codes, we penalize similarity between negative pairs and distance between positive pairs for each code. Finally, we incorporate these contrastive losses in a Variational Autoencoder (VAE) based architecture with VAE reconstruction loss and train the model in an end-to-end manner.
@@
With such a strategy, by just performing simple rotation to create contrastive pairs, DualContrast disentangles a much wider range of transformations caussing subtle changes in the pixel space,   e.g., viewpoint in Linemod, several writing styles in MNIST, etc. With DualContrast, we, for the first time, solved the scientific problem of disentangling protein identities from protein conformations (defined as subtle changes) in simulated 3D protein images (Fig. 5 and Fig. 6) inside the cell, known as cellular cryo-ET subtomograms. DualContrast could successfully disentangle the protein identities as content and protein conformations as transformation in simulated subtomograms, which is not achievable with previous methods. Overall, our qualitative and quantitative experimental results show the efficacy of DualContrast in isolating content from transformation compared to existing relevant methods.
@@
  • We develop a novel contrastive generative modeling strategy that creates positive-negative pairs for both content and transformation latent codes. We introduced a novel way of designing contrastive pairs for transformation from both data and latent space.
  • We show that DualContrast is effective in disentangling a wide range of transformations that causes subtle changes in pixel space, from the content in various shape focused image datasets.
  • Leveraging DualContrast, we, for the first time, disentangled protein identities (i.e. composition) from protein conformations in simulated 3D cellular cryo-ET subtomograms, as a proof of principle. DualContrast is the first method that can identify distinct conformations of proteins from protein mixture subtomogram datsets.
@@ 2 Related Works
Unsupervised Content-Transformation Disentanglement: There are a few [1, 4, 2, 15] approaches that deal with the unsupervised disentanglement of content and transformation. However, these methods perform explicit parameterization of the transformation codes as affine or some specific parameterized transformation. In contrast, our method does not impose any explicit parameterization on the transformation latent codes and does not face the issues of the explicit methods (Section 1).
@@
Contrastive Learning based Disentanglement: Contrastive learning is the primary building block of our method. Contrastive Learning with data augmentation with existing approaches, e.g. , InfoNCE loss-based SimCLR [16], etc., has been previously used by [15, 8] to isolate content from style. Though in several scenarios, style in shape focused images can be referred to as transformations, these works do not specifically focus on shape focused real-world images. Moreover, our experiments demonstrate that SimCLR with rotation augmentation in a standard encoder-decoder framework leads to very poor disentanglement of content and transformation (Table 1). Also, [15] only creates positive pairs w.r.t content and negative pair w.r.t. style with data augmentation in SimCLR. We have observed this scheme does not work well in our scenario. Unlike these works, our approach uses a novel strategy to create “negative” and “positive” pairs for both content and transformation of latent codes without any InfoNCE loss.
@@
Unsupervised Content-style disentanglement: Apart from [15], several methods [6, 17, 18, 19, 20, 21] exist that perform unsupervised content-style disentanglement, focusing on natural images. Unlike these methods, our work primarily focuses on disentangling content and transformation in shape focused image datasets. Moreover, we do not depend on any ImageNet pretrained models as our images of interest differ greatly from the natural images of ImageNet. The very recent work by [6] assumes access to the style factors to disentangle that style from content in feature outputs from pretrained models. Unlike this work, we do not assume access to the transformations beforehand for disentanglement.
@@
Protein Composition-Conformation Disentanglement: One of the major contributions of our work is disentangling protein identity characterized by large compositional variability as the content codes and protein states characterized by conformational variability or subtle compositional variability as transformation codes from 3D cellular cryo-electron tomography (cryo-ET) subimages or subtomograms. Previously, Harmony [4] was used to disentangle large compositional variability in cellular subtomograms. HEMNMA-3D [22] and TomoFlow [23] were used to disentangle subtle compositional or conformational variability in subtomograms using known templates. SpatialVAE [2] disentangled subtle conformational variability from 2D cryo-EM images. Ours is the first work to deal with both large and subtle compositional variability in two different latent spaces for 3D cryo-ET subtomograms.
Further discussions on the related works can be found in Appendix. Section A.1.
@@ 3.1 Disentangling Content and Transformation
  • Condition 1: $\forall$ $T\in\mathcal{T}$ and $\forall$ $\mathbf{x}\in\mathcal{X}$, $h_{c}(T(\mathbf{x}))=h_{c}(\mathbf{x})$.
  • Condition 2: $\exists$ $T\in\mathcal{T}$ and $\forall$ $\mathbf{x}^{(1)},\mathbf{x}^{(2)}\in\mathcal{X}$, $h_{z}(T(\mathbf{x}^{(1)}))=h_{z}(\mathbf{x}^{(2)})$.
@@
For content-transformation (c-z) disentangling, changes in $\mathcal{C}$ must be unaffected by changes in $\mathcal{Z}$ and vice versa. In our problem setting, no labels associated with content space $C$ or transformation space $Z$ are known. The exact family of transformation $\mathcal{T}$ is also not known beforehand. Thus, we aim to simultaneously learn the family of transformations $\mathcal{T}$ and the content factors unaltered by $\mathcal{T}$.
@@ 3.2 Method Overview and Notation
Our method is built upon a variational auto-encoder (VAE) based architecture (Fig. 2), where we perform inference of the content and transformation factors in the data. Consider, a sample space $\mathcal{X}$, two latent distribution spaces $\Phi_{Z}$ and $\Phi_{C}$, two latent spaces $\mathcal{Z}$ and $\mathcal{C}$. We apply an encoder network $q_{\psi}:\mathcal{X}\rightarrow\mathcal{C}\times\mathcal{Z}$ on a sample space $bx\in\mathcal{X}$ to encode content distribution $\phi_{c}\in\Phi_{c}$ and transformation distribution $\phi_{z}\in\Phi_{z}$, where $(\phi_{c},\phi_{z})=q_{\psi}(\mathbf{x})$. Latent parameters $\mathbf{c}\in\mathcal{C}$, $\mathbf{z}\in\mathcal{Z}$ are drawn from distributions $\phi_{c}$ and $\phi_{z}$ respectively. These parameters are then decoded using a decoder network $p_{\theta}:\mathcal{C}\times\mathcal{Z}\rightarrow\mathcal{X}$ to produce $x_{\text{recon}}=p_{\theta}(\mathbf{c},\mathbf{z})$.
3.3 Variational Inference of Content and Transformation Codes